# Vript: A Video Is Worth Thousands of Words

### v1

@@ Abstract
Advancements in multimodal learning, particularly in video understanding and generation, require high-quality video-text datasets for improved model performance. Vript addresses this issue with a meticulously annotated corpus of 12K high-resolution videos, offering detailed, dense, and script-like captions for over 420K clips. Each clip has a caption of ~145 words, which is over 10x longer than most video-text datasets. Unlike captions only documenting static content in previous datasets, we enhance video captioning to video scripting by documenting not just the content, but also the camera operations, which include the shot types (medium shot, close-up, etc) and camera movements (panning, tilting, etc). By utilizing the Vript, we explore three training paradigms of aligning more text with the video modality rather than clip-caption pairs. This results in Vriptor, a top-performing video captioning model among open-source models, comparable to GPT-4V in performance. Vriptor is also a powerful model capable of end-to-end generation of dense and detailed captions for long videos. Moreover, we introduce Vript-Hard, a benchmark consisting of three video understanding tasks that are more challenging than existing benchmarks: Vript-HAL is the first benchmark evaluating action and object hallucinations in video LLMs, Vript-RR combines reasoning with retrieval resolving question ambiguity in long-video QAs, and Vript-ERO is a new task to evaluate the temporal understanding of events in long videos rather than actions in short videos in previous works. All code, models, and datasets are available in https://github.com/mutonix/Vript.
@@ 1 Introduction
Dataset Domain Text Len Clips Duration Resolution
HowTo100M [21] Open 4.0 136M 134Kh 240p
ACAV100M [22] Open - 100M 278h -
HD-VILA-100M [23] Open 32.5 103M 371Kh 720p
WebVid-10M [13] Open ~12 10M ~52Kh 360p
MSVD [24] Open 8.7 1970 5.3h -
MSR-VTT [16] Open 9.3 10K 40h 240p
DiDeMo [25] Flickr 8.0 27K 87h -
ActivityNet [26] Action 13.5 100K 849h 144p-720p
YouCook2 [27] Cooking 8.8 14K 176h -
VATEX [28] Open 15.2 41K ~115h -
HD-VG-130M [6] Open ~10 130M ~180Kh 720p
Panda-70M [14] Open 13.2 70.8M 167Kh 720p
Vript (Ours) Open ~145 420K 1.3Kh 720p-2K
@@ 5 Vript-Hard
As multimodal models advance in performance, a more challenging benchmark is required to evaluate their capabilities. We propose a hard video understanding benchmark, dubbed Vript-Hard, consisting of three challenging tasks: HAL (Hallucination Evaluation), RR (Retrieval then Reasoning), and ERO (Event Re-ordering). We evaluate a large range of image LLMs, namely BLIP2 [34], InstructBLIP [35], Qwen-VL [36], LLaVA 1.6 34B [1], and video LLMs, namely VideoChatGPT [20], Video-LLaMA [30], VideoChat [29], VideoChat2 [7], ST-LLM [33]. We also evaluate sophisticated close-source models, namely Claude 3-Sonnet and Opus [37], GPT-4V [15]. More details about Vript-Hard can be checked in Appendix E.
@@ 5.3 Vript-ERO: A Temporal Understanding Benchmark of Long Videos
In Table 4, "-" means these models fail to give answers. Different from previous tasks that only have questions, Vript-ERO also contains long descriptions of scenes, which indicates these models are weak in processing long instructions. For models having scores, they only give the correct orders of all three scenes (@3) in about 20% of questions. In Figure 6 (b), we collect answers to the questions that are answered incorrectly and analyze the reasons. We observe that the models can be easily misled by the provided descriptions. For example, environment descriptions like sunlight may imply the morning or evening, however, these events may come from different days in the video rather than sequentially happening in one day. In 31.4% of cases, some events are absent in the input frames due to the limitation of the number of input images for models like GPT-4V. Besides, in 25.1% of cases, the models do not recognize which scene to be sequenced based on the descriptions.

### v2

@@ Abstract
Advancements in multimodal learning, particularly in video understanding and generation, require high-quality video-text datasets for improved model performance. Vript addresses this issue with a meticulously annotated corpus of 12K high-resolution videos, offering detailed, dense, and script-like captions for over 420K clips. Each clip has a caption of ~145 words, which is over 10x longer than most video-text datasets. Unlike captions only documenting static content in previous datasets, we enhance video captioning to video scripting by documenting not just the content, but also the camera operations, which include the shot types (medium shot, close-up, etc) and camera movements (panning, tilting, etc). By utilizing the Vript, we explore three training paradigms of aligning more text with the video modality rather than clip-caption pairs. This results in Vriptor, a top-performing video captioning model among open-source models, comparable to GPT-4V in performance. Vriptor is also a powerful model capable of end-to-end generation of dense and detailed captions for long videos. Moreover, we introduce Vript-Hard, a benchmark consisting of three video understanding tasks that are more challenging than existing benchmarks: Vript-HAL is the first benchmark evaluating action and object hallucinations in video LLMs, Vript-RR combines reasoning with retrieval resolving question ambiguity in long-video QAs, and Vript-ERO is a new task to evaluate the temporal understanding of events in long videos rather than actions in short videos in previous works. All code, models, and datasets are available in https://github.com/mutonix/Vript. PS: We have included more video-text datasets (Vript_CN & Vript_Multilingual) in the Vript series.
@@ 1 Introduction
Dataset Domain Text Len Clips Duration Resolution Lang
HowTo100M [21] Open 4.0 136M 134Kh 240p en
ACAV100M [22] Open - 100M 278h - en
HD-VILA-100M [23] Open 32.5 103M 371Kh 720p en
WebVid-10M [13] Open ~12 10M ~52Kh 360p en
YT-Temporal-180 [24] Open ~10 180M - 480p en
MSVD [25] Open 8.7 1970 5.3h - en
MSR-VTT [16] Open 9.3 10K 40h 240p en
DiDeMo [26] Flickr 8.0 27K 87h - en
ActivityNet [27] Action 13.5 100K 849h 144p-720p en
YouCook2 [28] Cooking 8.8 14K 176h - en
VATEX [29] Open 15.2 41K ~115h - en
HD-VG-130M [6] Open ~10 130M ~180Kh 720p en
Panda-70M [14] Open 13.2 70M 167Kh 720p en
InternVid [30] Open 17.6 234M 760.3Kh 720p en
Vript Open ~145 420K 1.3Kh 720p-2K en
Vript-CN Open ~150 293K - 720p-1080p zh
Vript-Multilingual Open ~150 677K - 720p-1080p multi
@@ 5 Vript-Hard
As multimodal models advance in performance, a more challenging benchmark is required to evaluate their capabilities. We propose a hard video understanding benchmark, dubbed Vript-Hard, consisting of three challenging tasks: HAL (Hallucination Evaluation), RR (Retrieval then Reasoning), and ERO (Event Re-ordering). We evaluate a large range of image LLMs, namely BLIP2 [36], InstructBLIP [37], Qwen-VL [38], LLaVA 1.6 34B [1], and video LLMs, namely VideoChatGPT [20], Video-LLaMA [32], VideoChat [31], VideoChat2 [7], ST-LLM [35], PLLaVA 7B [39], VILA-1.5 8B [40]. For open-source image and video LLMs, we sample 4 and 16 frames (following VideoChat2) per video respectively. We also evaluate sophisticated close-source models, namely Claude 3-Sonnet and Opus [41], GPT-4V [15], Claude 3.5-Sonnet [41], Gemini-1.5-Pro [42], GPT-4O [43]. For the close-source models, we sample 10 frames per video because GPT-4V can only accept a maximum of 10 frames. We also introduce a special setting for GPT-4O that we sample with 1 fps or a maximum of 100 frames (1 fps/100 fr). More details about Vript-Hard can be checked in Appendix E.
@@ 5.3 Vript-ERO: A Temporal Understanding Benchmark of Long Videos
In Table 4, "-" means these models fail to give answers. Different from previous tasks that only have questions, Vript-ERO also contains long descriptions of scenes, which indicates these models are weak in processing long instructions. For models having scores, they only give the correct orders of all three scenes (@3) in about 20% of questions. In Figure 6 (b), we collect answers to the questions that are answered incorrectly and analyze the reasons. We observe that the models can be easily misled by the provided descriptions. For example, environment descriptions like sunlight may imply the morning or evening, however, these events may come from different days in the video rather than sequentially happening in one day. In 31.4% of cases, some events are absent in the input frames due to the limitation of the number of input images for models like GPT-4V. Besides, in 25.1% of cases, the models do not recognize which scene to be sequenced based on the descriptions. For the GPT-4O (1 fps/100 fr), which operates at 1 fps or 100 frames, an increased number of frames within the input significantly enhances the overall scores. This is because the probability of the relevant events being omitted decreases with a larger input frame count.